How do you solve 2y=5x-1 and x=y+2? (Solve by graphing)
How do you solve x=5 and 3y=2x-1 (Solve by graphing)
How do you solve y=4 and 2x-y=1 (Solve by graphing)
How do you solve y-5x=10 and x-5y=10 (Solve by graphing)
i don't understand how to do these!:(

Answer by Alan3354(69443)

All of these are linear equations, so their graphs are straight lines.
A straight needs only 2 points to graph it.
----------------------
How do you solve 2y=5x-1 and x=y+2? (Solve by graphing)
2y = 5x-1
y = (5/2)x - 1/2
Pick 2 values of x and find y
x...y
1 2 --> (1,2)
3...7 --> (3,7)
Plot those 2 points and draw a line thru them.
Do the same for x=y+2
y = -x+2
Find 2 points, plot them draw the line.
Where the 2 lines intersect is the solution.
-------------------------
How do you solve x=5 and 3y=2x-1 (Solve by graphing)
This one is easier as x=5 is a vertical line thru (5,0).
3y=2x-1
y = (2/3)x - 1/3
Do it like the 1st one.
-------------------------
How do you solve y=4 and 2x-y=1 (Solve by graphing)
How do you solve y-5x=10 and x-5y=10 (Solve by graphing)
--------------
email me via the thank you note for help, or to check your work.